A boat can travel at a speed of 12km / h in still water. If it takes three hours to go 15km upstream and back, find the speed of the current.
Answer by stanbon(75887)

A boat can travel at a speed of 12km / h in still water. If it takes three hours to go 15km upstream and back, find the speed of the current.
-----------------------
Upstream DATA:
rate = 12-c km/h ; distance = 15 km ; time = 15/(12-c) hrs
------
Downstream DATA:
rate = 12+c km/h ; distance = 15 km ; time = 15/(12+c) hrs
--------
Equation:
time + time = 3 hrs.
15/(12-c) + 15/(12+c) = 3
-----
[15(12+c) + 15(12-c) = 3(144-c^2)
-----
360 = 3(144-c^2
120 = 144-c^2
c^2 = 24
c = 2sqrt(6) km/h (speed of the current)
==============
Cheers,
Stan H.
===============